OpenStudy (anonymous):

The following equation is true. How is the left simplified to match the right? This has been bugging me for hours now.

OpenStudy (anonymous):

\[\sqrt[3]{7+5 \sqrt{2} } = 1 + \sqrt{2}\]

OpenStudy (anonymous):

Try cubing both sides.

OpenStudy (anonymous):

\[\large 7+5\sqrt{2}=(1+\sqrt{2})^3\]

OpenStudy (anonymous):

Sorry I wasn't clear on this: Considering \[\sqrt[3] {7+5 \sqrt {2}}\] by it self. What are the steps necessary to simplify the equation into a rational number. in this case I cannot cube both sides.

OpenStudy (zugzwang):

It does help to cube both sides though :) Makes working backwards much clearer

OpenStudy (anonymous):

Ah, verifying an identity, yes that does make it more difficult.

OpenStudy (zugzwang):

To be clear, cubing both sides is only a sketch of it... Perhaps it would be simpler if you write \[7 + 5\sqrt{2}\] as \[1 + 3\sqrt{2} + 6 + 2\sqrt{2}\]

OpenStudy (zugzwang):

Such that \[\sqrt[3]{7+5\sqrt{2}}=\sqrt[3]{1 + 3\sqrt{2}+6+2\sqrt{2}}\] \[=\sqrt[3]{\left( 1 + \sqrt{2} \right)^{3}}\]

OpenStudy (anonymous):

That defiantly answered my question on how to try thinking when doing so. Didn't think of the interior by it self as a sort of a quadratic equation.

OpenStudy (zugzwang):

A little creativity and clever manipulation (?) could help a whole lot :D